OpenStudy (anonymous):

I need steps for this one too! PLEASE HELP! 24*log(3x)=60

OpenStudy (anonymous):

Ok, this is the same deal. So lets approach it the same way. This time I'll instruct and you perform

OpenStudy (anonymous):

To start we need to isolate the log(3x) factor. To do that we have to get rid of the other factor 24. So we divide both sides by 24

OpenStudy (anonymous):

What do we have now?

OpenStudy (anonymous):

Polpak - using the principals before I divide both sides by 24 and I end up with log 10 (3x) = 2.5

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Now we re-write the log in exponential form

OpenStudy (anonymous):

then I have to put in in the format as before with the logb (a)=k I get log10 (3x)=2.5

OpenStudy (anonymous):

jenn.. did you delete my response?

OpenStudy (anonymous):

I had posted the complete solution..

OpenStudy (anonymous):

helpingtutors - did not get your response. sorry, working it through with polpak thanks

OpenStudy (anonymous):

Right, so lets write the other half of that relation: b^k = a

OpenStudy (anonymous):

10^2.5=3x

OpenStudy (anonymous):

the b (base) in this case is 10. our 'a' is 3x, and our k is 2.5

OpenStudy (anonymous):

answer was 105.4092

OpenStudy (anonymous):

That's correct jenn

OpenStudy (anonymous):

Now to isolate the x, we divide by the coefficient in front (the 3)

OpenStudy (anonymous):

so when I calculate it I get 316.227766/3=105.409

OpenStudy (anonymous):

do I have it right?

OpenStudy (anonymous):

Nicely done

OpenStudy (anonymous):

Thanks! I am now understanding this very tedious process!

OpenStudy (anonymous):

It will get faster =)

OpenStudy (anonymous):

I appreciate the time to explain it to me! That's how I learn, step-by-step! Thanks again! I awarded you another medal!

OpenStudy (anonymous):

Anytime. You're a good student to work with.